OpenStudy (misssmartiez):

The mass of a freight train is 9.98 ⋅ 10^6 kilograms. The mass of an aircraft carrier is 7.3 ⋅ 10^7 kilograms. Approximately how many more kilograms is the mass of an aircraft carrier than the mass of a freight train? Show your work and write your answer in scientific notation.

OpenStudy (misssmartiez):

@jhonyy9

OpenStudy (misssmartiez):

@Hero @TheSmartOne

OpenStudy (misssmartiez):

@perl

OpenStudy (misssmartiez):

@whpalmer4

OpenStudy (misssmartiez):

I simply need help on how to put these SCIENTIFIC NOTATIONS in an expanded format.

OpenStudy (misssmartiez):

Then I would take it from there. Please help, I am behind.

jhonyy9 (jhonyy9):

just divide the mass of aircraft by the mass of freight train

jhonyy9 (jhonyy9):

7,3 *10^7 ----------- = (7,3/9,98) *10^(7-6) = ? 9,98 *10^6

jhonyy9 (jhonyy9):

= 0,73*10 so is by 7,3 time more the mass of an aircraft than a mass of a freight train hope helped

jhonyy9 (jhonyy9):

TSO my answer is false ?

jhonyy9 (jhonyy9):

ty